The concentration of cholesterol dissolved in chloroform is 0.615g per 10ml of solution.A portion of this solution in a 5cm polarimeter tube caused an observation of -1.7, calculate the specific rotation of cholesterol?
1
Expert's answer
2020-03-12T08:22:43-0400

0.615 g – 10 mL

      X g – 100 mL

C = X = 0.615g × 100mL / 10mL = 6.15g (in 100mL).

[α]D20 = 100 α / (l × C) = 100 × (-1.7) / (0.5dm × 6.15g) = -55.28 ≈ -55.3.